OpenStudy (anonymous):

The density of trees (measured in trees per square mile) in a particular forest can be modeled by the following function: p(x) = 500/((x+2)^3) The forest is rectangular shaped. It borders a straight road for 8 miles while extending 3 miles away from the road. The x in the density function is measured in miles away from the road. (a) Write a definite integral giving the total number of trees in the forest. Express your answer using the form below.

OpenStudy (mathteacher1729):

What part of this problem are you having trouble with? (writing the integral, solving it?)

OpenStudy (anonymous):

in writing the integral

OpenStudy (mathteacher1729):

This problem statement is sort of wonky. There are a few ways to interpret it. I'd ask your teacher for clarification.

OpenStudy (anonymous):

I know that the integral is 0 to 3

OpenStudy (anonymous):

the integral is Density X Area

OpenStudy (anonymous):

so you can't help me? thank you so much

OpenStudy (mathteacher1729):

Sorry, phone rang. I think it's a strangely worded problem. Here is one way to interpret it: This gives us the number of trees in a 3 mile slice of one mile wide. \[\huge 1\text{ mile} \times \int_{x=0}^{x=3}\frac{500}{(x+2)^2}dx\] The forest is 8 miles wide... That should be enough hint for now and the answer I get is a whole number, so I'm fairly certain this is how it's supposed to be interpreted.

OpenStudy (anonymous):

why 500/((x+2)^2) and not 500/((x+2)^3)?

OpenStudy (anonymous):

I think you are wrong?

OpenStudy (anonymous):

@mathteacher1729 you are wrong

OpenStudy (mathteacher1729):

@Ldaniel because I pressed the 2 button instead of the 3 button. Sorry. :( Everything else about the interpretation seems to check out though.

OpenStudy (anonymous):

no it's wrong it not 500/((x+2)^3)

OpenStudy (anonymous):

it should be 500/((x+2)^3) X the Area. but i don't know how to find the area of of the forest

OpenStudy (anonymous):

never-mind I got it . the total number of trees is its 500/((x+2)^3) X 8 miles X deltal x

OpenStudy (mathteacher1729):

The area of the forest is the area of the rectangle. 3 miles x 8 miles = 24 square miles.

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\frac{ 500 }{ (x+2)^3 }\times 8 dx\]

OpenStudy (anonymous):

that's the right integral.

OpenStudy (mathteacher1729):

Bingo.

OpenStudy (anonymous):

you need to practice your calculus mathteacher :)

OpenStudy (anonymous):

thanks anyways :) you helped me a lot

OpenStudy (mathteacher1729):

Glad it was helpful, sorry for the typo where I wrote 2 instead of 3. The idea of my post was still correct. I gave you a 1 x 3 rectangle. You need an 8 x 3 rectangle. Multiply by 8 and you get your integral. :) nighty night.

OpenStudy (anonymous):

yeah thanks